Program has 110 participants
78 participants took CM 400
63 participants took CM 450
54 participants took both CM 400 and CM 450
How many participants took CM 400 alone?
How many participants took CM 450 alone?
How many participants did not take neither CM400 nor CM450?
How many students did not take CM 400?
How many students did not take CM 450?

Answer by mizzmake2(3)

Program has 110 participants
78 participants took CM 400
63 participants took CM 450
54 participants took both CM 400 and CM 450

How many participants took CM 400 alone?
78 participants
How many participants took CM 450 alone?
63 participants
How many participants did not take neither CM400 nor CM450?
110-54=
56 participants

How many students did not take CM 400?
110-78=
32 students

How many students did not take CM 450?
110-63=
47 students


This is not a for sure answer but this is how I would answer it...
Good Luck!!!